4. The total frictional and windage loss of a 3-phase induction motor is 300 W. With a 2% slip, find the %efficiency of the motor when it draws 25 A current at 0.8 p.f. lagging from a 460-V supply and the equivalent rotor resistance referred to stator Rgap = 8.0 ohm
Answer #1

Solution:

Given that WF AL = 300 w S : 0.02 Is = 85 A @ 0.8 log Vs = 460v Polp = √3 vs Is cos o rox460 X 25 X0.8 = 15934-86 .Pm = (1-6)